OpenStudy (anonymous):

Find a formula for (f(g(x)) assuming f(x)=3^(3+7x) and g(x)=log3 (x). I have tried it so many times but can't get the correct answer!

OpenStudy (anonymous):

\[f(x)=3^{3+7x}\\ g(x)=\log_3x\] \[\large f(g(x))=f(\log_3x)=3^{3+7\log_3x}=3^3\cdot3^{7\log_3x}\] Use the following properties: \[a\log_bx=\log_b(x^a)\\ b^{\log_bx}=x\]

OpenStudy (anonymous):

Thank you so much! Is that the simplest form it can be in?

OpenStudy (anonymous):

Have you applied the properties yet?

OpenStudy (anonymous):

I got it down to 27(3^log3x)^7 so 27x^7. Correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Sweet! Thank you for your help!